OpenStudy (anonymous):

AB is tangent to O if AO=24 and BC=50 what AB

OpenStudy (anonymous):

Since AB is tangent to O, and AO is a radius of the circle, then we know that AB is perpendicular to AO So we have a right triangle ABO Which allows us to use the pythagorean theorem a^2+b^2 = c^2 Plug in the given side lengths to get AO^2+AB^2 = BO^2 From there, plug in the given lengths AB = x, AO = 24, BO = 27+24 = 51 (note: BO = BC+CO, BC = 27 and CO = AO = 24) So this means... AO^2+AB^2 = BO^2 24^2+x^2 = 51^2 From here, solve for x.

OpenStudy (anonymous):

24^2+x^2 = 51^2 576+x^2 = 2601 x^2 = 2601-576 x^2 = 2025 x = sqrt(2025) x = 45 So AB = 45